OpenStudy (anonymous):

So i've completely forgot how to do fractions from last semester :(. -2/5y= 1/7 solve for y. Do i find the lcd again or?

OpenStudy (anonymous):

forgotten*

OpenStudy (whpalmer4):

\[\frac{-2y}{5}=\frac{1}{7}\]I would multiply by 35 (5*7) and solve. With only 1 term on each side, you could also do this by cross-multiplication: \[\frac{a}{b}=\frac{c}{d}\]\[ad=bc\]

OpenStudy (whpalmer4):

\[-2y*7=1*5\]\[-14y=5\]

OpenStudy (whpalmer4):

Sometimes that's easier than finding the common denominator if you have some hairy expressions in the fractions

OpenStudy (anonymous):

whpalmer is right. You can think of the fraction on the right becoming a "1", therefore the y is being isolated by itself. This is done by taking it's reciprocal. - 2/5 times -5/2 is one. Right? Whatever you do on one side an equal side you must do it on the other.

OpenStudy (anonymous):

Excuse me, you can think of the fraction on the left* :)